OpenStudy (anonymous):

what is the distance from the orgin to the midpoint of the line segment defined by the endpoints (=4,0) and (2,-2)

OpenStudy (anonymous):

hi agian

OpenStudy (anonymous):

first of all find the mid point

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

\[x_m=\frac{ x_1+x_2 }{ 2 }\]

OpenStudy (anonymous):

\[y_m=\frac{ y_1+y_2 }{ 2 }\]

OpenStudy (anonymous):

x=(-4+2)/2=-1

OpenStudy (anonymous):

-3,-1

OpenStudy (anonymous):

is the answer i got at the end

OpenStudy (anonymous):

wait u have an error in the question

OpenStudy (anonymous):

do u mean -4?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok then take the distance to the origin which (0,0)

OpenStudy (anonymous):

do u know the law of distance?

OpenStudy (anonymous):

the formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the root one

OpenStudy (anonymous):

d=sqrt[(x2-x1)^2 + (y2-y1)^2]

OpenStudy (anonymous):

d=sqrt(10)

OpenStudy (anonymous):

i get 36, -4

OpenStudy (anonymous):

listen ur mid point is -3,1 right

OpenStudy (anonymous):

9+1 under the root

OpenStudy (anonymous):

from origin

OpenStudy (anonymous):

unless u calculated the mid point in a wrong way

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

?